A cheetah can run at a maximum speed
92 km/h and a gazelle can run at a maximum speed of 72.7 km/h.
If both animals are running at full speed,
with the gazelle 72.9 m ahead, how long before
the cheetah catches its prey?
Answer in units of s. Don't round answer.
1
Expert's answer
2019-10-10T09:53:33-0400

Lets choose coordinate system so that cheetah is at the origin of "x"-axis and gazelle "L" meters ahead.

The equations of motion of cheetah and gazelle are "x_C(t) = v_C t", "x_G(t) = v_G t + L" respectively, where "v_C = 92 \frac{km}{h}" , "v_G = 72.7 \frac{km}{h}", "L = 72.9 m". When cheetah catches the gazelle, "x_C(t) = x_G(t)", from where "v_C t = v_G t + L \Rightarrow t = \frac{L}{v_C - v_G} \approx 13.6 s".